A proton and an electron, both at rest initially, combine to form a hydrogen atom in the ground state. A single photon is emitted in this process. What is its wavelength?


1
Expert's answer
2021-11-24T19:12:56-0500

The wavelength is


"\lambda=\frac{hc}E=\frac{hc}{\Delta mc^2}=\frac{h}{\Delta mc}=67.9\text{ pm}."

In this equation:


"\Delta m=m_\text{H}-(m_p+m_e)."